# bibliography.bib

### v1

@@ Abstract
For the group GL(n), we construct an action of the derived category of coherent sheaves on the Grothendieck-Springer resolution on a certain subcategory of a finite monodromic Hecke category. We use this to construct a partial categorification of the projection from the extened affine to the finite Hecke algebra of GL(n). As a crucial intermediate step, we compute the exterior powers, with respect to the perversely truncated multiplicative convolution, of a parabolic Springer sheaf corresponding to a maximal parabolic subgroup fixing a line in the defining $n$-dimensional representation of GL(n).
@@
For a stack $X$ defined over $\mathbf{k}$, let $D^{b}(X)$ be the bounded $\ell$-adic derived category of $\overline{\mathbf{Q}}_{\ell}$-sheaves with constructible cohomology. Let $P(X)$ be the abelian category of perverse sheaves on $X$ with respect to the middle perversity. For a group $H$ acting on a scheme $X$, let $D^{b}_{H}(X)=D^{b}(X/H)$, $P_{H}(X)$ be the corresponding equivariant categories, defined as in [bernsteinEquivariantSheavesFunctors2006], [laszloSixOperationsSheaves2008a]. Let $\underline{\overline{\mathbf{Q}}_{\ell}}_{X}$ stand for the constant sheaf on $X$.
For an arbitrary stack $\mathbf{X}$ over $\mathbf{K}$, write $\mathrm{coh}(\mathbf{X}),D^{b}_{\mathrm{coh}}(\mathbf{X})$ for the category of coherent sheaves and bounded coherent derived category of sheaves on $\mathbf{X}$, respectively, and let $\mathrm{perf}(\mathbf{X})\subset D^{b}_{\mathrm{coh}}(\mathbf{X})$ stand for the perfect subcategory. For a group $\mathbf{H}$ acting on a scheme $\mathbf{X}$, let $D^{b}_{\mathrm{coh},\mathbf{H}}(\mathbf{X})=D^{b}_{\mathrm{coh}}(\mathbf{X}/\mathbf{H})$, $\mathrm{perf}_{\mathbf{H}}(\mathbf{X})$ be the corresponding equivariant categories.
@@
We write $\operatorname{Rep}(\mathbf{G})$ for the symmetric monoidal category of finite dimensional representations of $\mathbf{G}$ defined over $\mathbf{K}$, $\operatorname{Rep}(\mathbf{G})\simeq\mathrm{coh}(\operatorname{Spec}(\mathbf{K})/\mathbf{G})$.
For $\mathbf{G}=\mathbf{GL}_{n}$ we identify the weight lattice $\Lambda$ of $\mathbf{G}$ with $\mathbb{Z}^{n}$ with basis $\mathbf{e}_{1},\dots,\mathbf{e}_{n}$, so that the Weyl group $W=S_{n}$ of $\mathbf{GL}_{n}$ acts by permutations of this basis, and simple roots are given by $\alpha_{i}=\mathbf{e}_{i+1}-\mathbf{e}_{i},i=1,\dots,n-1$. Write $s_{i}=s_{\alpha_{i}}\in W$. We adopt similar conventions for $G=\textrm{GL}_{n}$, defining a basis $e_{1},\dots,e_{n}$. Let $w_{0}$ stand for the longest element in $W$.
@@
Let $\mathcal{T}$ be a $\mathbf{K}$-linear triangulated category, and let $X$ be a stack defined over $\mathbf{K}$. We call an exact action of the monoidal category $\mathrm{perf}(\mathbf{X})$ on $\mathcal{T}$ a linear structure on $\mathcal{T}$ over $X$
@@
Let $\mathcal{H}(G)$ stand for the completed unipotently monodromic Hecke category of a reductive group $G$, defined in [bezrukavnikovKoszulDualityKacMoody2014]. This is a certain completion of a full subcategory of unipotently monodromic complexes in $D^{b}(U\backslash G/U)$. $\mathcal{H}(G)$ is a monoidal category, monoidal structure being given by the convolution operation, which we denote by $\star$-- ⋆ -.
Let $\mathfrak{g}$ be the Lie algebra of $\mathbf{G}$, $\mathfrak{b}$ be the Lie algebra of a Borel subgroup $\mathbf{B}\subset\mathbf{G}$, and let $\tilde{\mathfrak{g}}=\mathbf{G}\times^{B}\mathfrak{b}$ be the Grothendieck-Springer resolution. The goal of this paper is to prove that the Hecke category $\mathcal{H}_{n}=\mathcal{H}(\textrm{GL}_{n})$ associated to the general linear group $G=\textrm{GL}_{n}$ admits the linear structure over the adjoint quotient of the Steinberg variety $\mathbf{St}_{\mathfrak{g}}=\tilde{\mathfrak{g}}\times_{\mathfrak{g}}\tilde{\mathfrak{g}}$ associated to the Lie algebra $\mathfrak{g}=\mathfrak{gl}_{n}$. For the rest of the introduction, let $G=\textrm{GL}_{n},\mathbf{G}=\mathbf{GL}_{n}$.
The following theorem is one of our main results, see Theorem 4.1.1.
@@
where the monoidal structure on the perfect equivariant category $\mathrm{perf}_{G}(\mathbf{St}_{\mathfrak{g}})$ is given by convolution.
The image of this functor, denoted by $\mathcal{H}_{n}^{\mathrm{perf}}$, is the full triangulated subcategory of $\mathcal{H}_{n}$ generated by the big pro-unipotent tilting sheaf $\hat{\mathcal{T}}_{w_{0}}$, see Section 2. To prove the above theorem, we equip $\mathcal{H}_{n}^{\mathrm{perf}}$ with two linear structures over the stack $\tilde{\mathfrak{g}}$.
@@
We proceed to outline the main ideas of our construction of the linear structures and the functor $\varpi$. The construction employs the Tannakian formalism for $D^{b}_{\mathrm{coh}}(\tilde{\mathfrak{g}}/\mathbf{G}),\mathrm{perf}_{\mathbf{G}}(\mathbf{St}_{\mathfrak{g}})$ developed by Bezrukavnikov in [bezrukavnikovTwoGeometricRealizations2016]. Roughly, this formalism allows to construct a linear structure over $\tilde{\mathfrak{g}}/G$ on an arbitrary monoidal triangulated category $\mathcal{T}$ by equipping $\mathcal{T}$ with an action of $\operatorname{Rep}(\mathbf{G}\times\mathbf{T})$ satisfying certain compatibilities. If we have two such structures with isomorphic actions of two copies of $\operatorname{Rep}(\mathbf{G})$, we get a monoidal functor $\mathrm{perf}_{\mathbf{G}}(\mathbf{St}_{\mathfrak{g}})\to\mathcal{T}$.
@@
In [bezrukavnikovTwoGeometricRealizations2016], this formalism is used to construct a monoidal equivalence between $D^{b}_{\mathrm{coh},\mathbf{G}}(\widehat{\mathbf{St}_{\mathfrak{g}}})$, where $\widehat{\mathbf{St}_{\mathfrak{g}}}$ stands for a certain formal completion of $\mathbf{St}_{\mathfrak{g}}$, and the completed monodromic affine Hecke category $\mathcal{H}_{G^{\vee}}^{\mathrm{aff}}$. We do not know if it is possible to extend the functor $\varpi$ to $D^{b}_{\mathrm{coh},\mathbf{G}}(\widehat{\mathbf{St}_{\mathfrak{g}}})$.
@@
The functor $\varpi$ constructed in this paper partially categorifies a homorphism from the affine to the finite Hecke algebra of $\textrm{GL}_{n}$.
The homomorphism in question may be most readily seen on the level of the corresponding braid groups. If we interpret the braid group $B_{n}$ on $n$ strands as the fundamental group of the configuration space $\textrm{Conf}_{n}(\mathbb{C})$ of $n$ points in the complex plane $\mathbb{C}$ based at a configuration avoiding 00, and the affine braid group ${B}_{n}^{\textrm{aff}}$ as the fundamental group of the configuration space $\textrm{Conf}_{n}(\mathbb{C}^{*})$ of $n$ points in the punctured complex plane $\mathbb{C}^{*}$ based at the same configuration, then the inclusion $\mathbb{C}^{*}\to\mathbb{C}$ gives a homomorpism
@@
between the corresponding extended affine Hecke algebra ${H}^{\textrm{aff}}_{n}$ of the reductive group $\mathbf{GL}_{n}$ and the finite Hecke algebra $H_{n}$. Recall, that ${H}^{\textrm{aff}}_{n}$ contains $H_{n}$ as a subalgebra. It also contains a commutative subalgebra, defined by Bernstein, which is identified with the group algebra of the (co-)weight lattice of $\mathbf{GL}_{n}$. Let $\theta_{1},\dots,\theta_{n}$ be the generators of this subalgebra corresponding to the basis elements $\mathbf{e}_{1},\dots,\mathbf{e}_{n}$. The homomorphism $\varpi_{\textrm{alg}}$ is defined uniquely by the following two properties: $\varpi_{\textrm{alg}}(x)=x$ for $x\in H_{n}$; $\varpi_{\textrm{alg}}(\theta_{1})=1.$ The images of the elements $\theta_{k}$ in the finite Hecke algebra are known as multiplicative Jucys-Murphy elements therein, and play an important role in the representation theory of $H_{n}$, analogous to the role of classical Jucys-Murphy elements for symmetric groups, see [isaevRepresentationsHeckeAlgebras2005], [okounkovNewApproachRepresentation1996].
$\operatorname{Rep}(\mathbf{T})$-action employed in the Tannakian formalism of [bezrukavnikovTwoGeometricRealizations2016] described above categorifies the action of the commutative subalgebra generated by $\theta_{i}^{\pm 1}$. The action of $V\in\operatorname{Rep}(\mathbf{G})$ categorifies the action of the central element in $H^{\textrm{aff}}_{n}$ corresponding to $V$, see [lusztigSingularitiesCharacterFormulas1983], [gaitsgoryConstructionCentralElements2001]. This element can be expressed as a symmetric Laurent polynomial in $\theta_{i}$, according to the character of $V$, which is expressed on the categorical level as the compatibility condition on $\operatorname{Rep}(\mathbf{T})$- and $\operatorname{Rep}(\mathbf{G})$- actions mentioned above.
@@
The action of $\operatorname{Rep}(\mathbf{T})$ is defined in Section 2.2 using the description of the homomorphism $\varpi_{\textrm{br}}$ and the braid group action on $\mathcal{H}_{n}$ defined by Rouquier [rouquierCategorificationBraidGroups2004]. To any braid $\beta\in B_{n}$ one can attach an object $F_{\beta}\in\mathcal{H}_{n}$. For any weight $\lambda$ there is a lift $\tilde{\theta}(\lambda)\in B_{n}^{\textrm{aff}}$ of the Bernstein commutative subalgebra element $\theta(\lambda)\in H_{n}^{\textrm{aff}}$ attached to $\lambda$. We define the action of $\lambda$, considered as an object in $\operatorname{Rep}(\mathbf{T})$, to be the functor of convolution with $F_{\varpi_{\textrm{br}}(\tilde{\theta}(\lambda))}$. This complex can be expressed as a convolution of complexes of the form $F_{\beta}$ attached to the Jucys-Murphy braids.
@@
The operation $\star$ is sometimes called *-convolution, to distinguish it from the dual operation using $m_{!}$ instead of $m_{*}$ in its definition. In this paper, we will work with the *-convolution $\star$-- ⋆ - and refer to it simply as convolution. This operation is a categorification of the operation of convolution of functions on $G$.
The equivariant category $D^{b}_{G}(G)$ can be equipped with the compatible monoidal structure. The resulting monoidal category is canonically braided. In particular, for any pair of objects $\mathcal{F},\mathcal{G}\in D^{b}_{G}(G)$, there is a braiding isomorphism
@@
$\theta:\operatorname{Id}_{D^{b}_{G}(G)}\to\operatorname{Id}_{D^{b}_{G}(G)},\theta_{\mathcal{F}}:\mathcal{F}\to\mathcal{F},\mathcal{F}\in D^{b}_{G}(G),$
@@
See [boyarchenkoCharacterSheavesUnipotent2013] for a detailed discussion.
The convolution $\star$ on $D^{b}(G)$ is right t-exact with respect to the perverse $t$-structure, since $m$ is an affine morphism. This allows to equip $P(G),P_{G}(G)$ with a perversely truncated convolution
@@
This makes $P_{G}(G)$ a braided monoidal category as well.
The canonical braided structure on $D^{b}_{G}(G)$ is almost never symmetric, even in the case when $G$ is commutative.
Let $G=\textrm{GL}_{n}$. It is well known that $G$ satisfies the following property: for any $g\in G$, the centralizer $Z_{G}(g)$ is connected. One checks, see Section 3.2, that in this case $P_{G}(G)$ with the monoidal structure $\star^{0}$ is symmetric braided. This essentially follows from the fact that the twist automorphism $\theta$ acts as identity on perverse objects. Consequently, we can speak of various Schur functors in this category, in particular of exterior powers $\wedge^{k}\mathcal{F}$ for any object $\mathcal{F}\in P_{G}(G)$. In this paper, we compute the exterior powers of a particular object called the parabolic Springer sheaf (with respect to a particular parabolic subgroup).
@@
Let $\pi:\tilde{\mathcal{N}}_{P}\to G$ be the natural projection, and define $\mathcal{S}_{P}=\pi_{*}\underline{\overline{\mathbf{Q}}_{\ell}}_{\tilde{\mathcal{N}}_{P}}[2\dim U].$
Let $\mathcal{N}_{G}$ stand for the variety of unipotent elements in $G$. Sheaves $\mathcal{S}_{P}$ are supported on $\mathcal{N}_{G}.$
Let now $G=\textrm{GL}_{n}=\textrm{GL}(V)$, for a fixed $n$-dimensional $\mathbf{k}$-vector space $V$, let $B$ be a Borel subgroup, and let $P_{1}\supset B$ be the parabolic subgroup fixing a line in $\mathbb{P}(V)$. Write
@@
Let $W$ stand for the Weyl group of $G$. It is isomorphic to the symmetric group $S_{n}$. Let
$\operatorname{Rep}(W)\to D^{b}_{G}(\mathcal{N}_{G}),V\mapsto\mathcal{S}_{V}$
stand for the functor given by the Springer correspondence, see [brylinskiTransformationsCanoniquesDualite1986], [borhoRepresentationsGroupesWeyl1981]. We follow the conventions of [brylinskiTransformationsCanoniquesDualite1986], so that the trivial representation of $W$ corresponds to the punctual perverse sheaf at the unit $e\in\mathcal{N}_{G}$. Consider a representation of $W$ on $V$ by permutations of a basis. We have $\mathcal{S}_{n}\simeq\mathcal{S}_{V}$. Write $\mathcal{S}_{\wedge^{k}}:=\mathcal{S}_{\wedge^{k}V}$.
Our main result is the following
@@
Exterior powers of $\mathcal{S}_{n}$ with respect to the symmetric monoidal structure $\star^{0}$ satisfy the following:
  1. a)
    $\wedge^{k}\mathcal{S}_{n}\simeq\mathcal{S}_{\wedge^{k}},$ for $k=0,\dots,n$.
  2. b)
    $\wedge^{k}\mathcal{S}_{n}\simeq 0$ for $k>n$.
It is well-known that the category $D^{b}_{G}(G)$, where $G$ acts on itself by adjoint action, acts on $\mathcal{H}_{n}$ by left and right convolution. Moreover, this action is central, meaning that the action on the right can be identified with the action on the left, using the equivariance with respect to adjoint action. It is thus natural to look for a $\operatorname{Rep}(\mathbf{G})$-action on $\mathcal{H}_{n}$ that factors through the action of $D^{b}_{G}(G)$.
Let $\operatorname{Rep}(\mathbf{G})^{+}$ stand for the symmetric monoidal category of polynomial representations of $G=\textrm{GL}_{n}$. We have the following corollary of Theorem 1.3.1, see Section 3.5.1 and the proof of Corollary 3.5.2.
@@
There is a monoidal functor $\operatorname{Rep}(\mathbf{G})^{+}\to P_{G}(G)$.
@@
1.3.3   Jucys-Murphy filtrations and monodromy.
To complete the construction of the linear structure and the functor $\varpi$ one needs to establish the compatibility of $\operatorname{Rep}(\mathbf{G})$ and $\operatorname{Rep}(\mathbf{T})$ actions. This is done in Section 4. The main idea is to use that the objects $\wedge^{k}\mathcal{S}_{n}$ come frome the parabolic induction from the Levi subgroup $\textrm{GL}_{k}\boxtimes\textrm{GL}_{n-k}$ of $G$. This allows to construct filtrations of the functors defining the $\operatorname{Rep}(\mathbf{G})$-action by the functors defining the $\operatorname{Rep}(\mathbf{T})$-action categorifying the weight filtration. The monodromy transformations required by the Tannakian formalism of [bezrukavnikovTwoGeometricRealizations2016] comes from the monodromy along the central $\mathbb{G}_{m}\subset\textrm{GL}_{k}$.

### v2

# Linear structure on a finite Hecke category in type A

@@ Abstract
For the group $\textrm{GL}_{n}$, we construct an action of the equivariant derived category of coherent sheaves on the Grothendieck-Springer resolution on a certain subcategory of a finite monodromic Hecke category. We use this to construct a partial categorification of the projection from the extened affine to the finite Hecke algebra of $\textrm{GL}_{n}$. As a crucial intermediate step, we compute the exterior powers, with respect to the perversely truncated multiplicative convolution, of a parabolic Springer sheaf corresponding to a maximal parabolic subgroup fixing a line in the defining $n$-dimensional representation of $\textrm{GL}_{n}$.
@@
For a stack $X$ defined over $\mathbf{k}$, let $D^{b}(X)$ be the bounded $\ell$-adic derived category of $\overline{\mathbf{Q}}_{\ell}$-sheaves with constructible cohomology. Let $P(X)$ be the abelian category of perverse sheaves on $X$ with respect to the middle perversity. For a group $H$ acting on a scheme $X$, let $D^{b}_{H}(X)=D^{b}(X/H)$ and $P_{H}(X)$ be the corresponding equivariant categories, defined as in [BL06], [LO08]. Let $\underline{\overline{\mathbf{Q}}_{\ell}}_{X}$ stand for the constant sheaf on $X$.
For an arbitrary stack $\mathbf{X}$ over $\mathbf{K}$, write $\mathrm{coh}(\mathbf{X})$ and $D^{b}_{\mathrm{coh}}(\mathbf{X})$ for the category of coherent sheaves and bounded coherent derived category of sheaves on $\mathbf{X}$, respectively, and let $\mathrm{perf}(\mathbf{X})\subset D^{b}_{\mathrm{coh}}(\mathbf{X})$ stand for the perfect subcategory. For a group $\mathbf{H}$ acting on a scheme $\mathbf{X}$, let $D^{b}_{\mathrm{coh},\mathbf{H}}(\mathbf{X})=D^{b}_{\mathrm{coh}}(\mathbf{X}/\mathbf{H})$ and $\mathrm{perf}_{\mathbf{H}}(\mathbf{X})$ be the corresponding equivariant categories.
@@
We write $\operatorname{Rep}(\mathbf{G})$ for the symmetric monoidal category of finite dimensional representations of $\mathbf{G}$ defined over $\mathbf{K}$, so that $\operatorname{Rep}(\mathbf{G})\simeq\mathrm{coh}(\operatorname{Spec}(\mathbf{K})/\mathbf{G})$.
For $\mathbf{G}=\mathbf{GL}_{n}$ we identify the weight lattice $\Lambda$ of $\mathbf{G}$ with $\mathbb{Z}^{n}$ with basis $\mathbf{e}_{1},\dots,\mathbf{e}_{n}$, so that the Weyl group $W=S_{n}$ of $\mathbf{GL}_{n}$ acts by permutations of this basis, and simple roots are given by $\alpha_{i}=\mathbf{e}_{i+1}-\mathbf{e}_{i}$ for $i=1,\dots,n-1$. Write $s_{i}=s_{\alpha_{i}}\in W$. We adopt similar conventions for $G=\textrm{GL}_{n}$, defining a basis $e_{1},\dots,e_{n}$. Let $w_{0}$ stand for the longest element in $W$.
@@
Let $\mathcal{T}$ be a $\mathbf{K}$-linear triangulated category, and let $X$ be a stack defined over $\mathbf{K}$. We call an exact action of the monoidal category $\mathrm{perf}(\mathbf{X})$ on $\mathcal{T}$ a linear structure on $\mathcal{T}$ over $X$ .
@@
[…]
Let $\mathfrak{g}$ be the Lie algebra of $\mathbf{G}$, let $\mathfrak{b}$ be the Lie algebra of a Borel subgroup $\mathbf{B}\subset\mathbf{G}$, and let $\tilde{\mathfrak{g}}=\mathbf{G}\times^{B}\mathfrak{b}$ be the Grothendieck-Springer resolution. The goal of this paper is to prove that the Hecke category $\mathcal{H}_{n}=\mathcal{H}(\textrm{GL}_{n})$ associated with the general linear group $G=\textrm{GL}_{n}$ admits a linear structure over $\mathbf{St}_{\mathfrak{g}}/\mathbf{G}$, where $\mathbf{G}=\mathbf{GL}_{n},$ and $\mathbf{St}_{\mathfrak{g}}=\tilde{\mathfrak{g}}\times_{\mathfrak{g}}\tilde{\mathfrak{g}}$ is the Steinberg variety associated with the Lie algebra $\mathfrak{g}=\mathfrak{gl}_{n}$. For the rest of the introduction, let $G=\textrm{GL}_{n},\mathbf{G}=\mathbf{GL}_{n}$.
Let $\mathcal{H}_{n}^{\mathrm{perf}}$ be the full triangulated subcategory of $\mathcal{H}_{n}$ generated by the big free-monodromic tilting sheaf $\hat{\mathcal{T}}_{w_{0}}$ – see Section 2. We will equip $\mathcal{H}_{n}^{\mathrm{perf}}$ with two compatible linear structures over the stack $\tilde{\mathfrak{g}}/\mathbf{G}$.
The following theorem is one of our main results (Theorem 4.1.1 in the main text).
@@
where the monoidal structure on the perfect equivariant category $\mathrm{perf}_{G}(\mathbf{St}_{\mathfrak{g}})$ is given by convolution. The functor takes values in $\mathcal{H}_{n}^{\mathrm{perf}}$ and is compatible with the linear structures over $\tilde{\mathfrak{g}}/\mathbf{G}$.
@@
We proceed to outline the main ideas of our construction of the linear structures and the functor $\varpi$. The construction employs the Tannakian formalism for $D^{b}_{\mathrm{coh}}(\tilde{\mathfrak{g}}/\mathbf{G}),\mathrm{perf}_{\mathbf{G}}(\mathbf{St}_{\mathfrak{g}})$ developed by Bezrukavnikov in [Bez16]. Roughly, this formalism allows us to construct a linear structure over $\tilde{\mathfrak{g}}/G$ on an arbitrary monoidal triangulated category $\mathcal{T}$ by equipping $\mathcal{T}$ with an action of $\operatorname{Rep}(\mathbf{G}\times\mathbf{T})$ satisfying certain compatibilities. If we have two such structures with isomorphic actions of two copies of $\operatorname{Rep}(\mathbf{G})$, we get a monoidal functor $\mathrm{perf}_{\mathbf{G}}(\mathbf{St}_{\mathfrak{g}})\to\mathcal{T}$.
@@
[…]
@@
Construction of the functor $\varpi$ is a step towards a categorification of a homomorphism from the affine to the finite Hecke algebra of $\textrm{GL}_{n}$.
The homomorphism in question may be most readily seen on the level of the corresponding braid groups. If we interpret the braid group $B_{n}$ on $n$ strands as the fundamental group of the configuration space $\textrm{Conf}_{n}(\mathbb{C})$ of $n$ points in the complex plane $\mathbb{C}$ based at a configuration avoiding 0, and the affine braid group ${B}_{n}^{\textrm{aff}}$ as the fundamental group of the configuration space $\textrm{Conf}_{n}(\mathbb{C}^{*})$ of $n$ points in the punctured complex plane $\mathbb{C}^{*}$ based at the same configuration, then the inclusion $\mathbb{C}^{*}\to\mathbb{C}$ gives a homomorpism
@@
between the corresponding extended affine Hecke algebra ${H}^{\textrm{aff}}_{n}$ of the reductive group $\mathbf{GL}_{n}$ and the finite Hecke algebra $H_{n}$. Recall that ${H}^{\textrm{aff}}_{n}$ contains $H_{n}$ as a subalgebra. It also contains a commutative subalgebra, defined by Bernstein, which is identified with the group algebra of the (co-)weight lattice of $\mathbf{GL}_{n}$. Let $\theta_{1},\dots,\theta_{n}$ be the generators of this subalgebra corresponding to the basis elements $\mathbf{e}_{1},\dots,\mathbf{e}_{n}$. The homomorphism $\varpi_{\textrm{alg}}$ is defined uniquely by the following two properties: $\varpi_{\textrm{alg}}(x)=x$ for $x\in H_{n}$, and $\varpi_{\textrm{alg}}(\theta_{1})=1.$ The images of the elements $\theta_{k}$ in the finite Hecke algebra are known as multiplicative Jucys–Murphy elements therein, and play an important role in the representation theory of $H_{n}$, analogous to the role of classical Jucys–Murphy elements for symmetric groups – see [IO05], [OV96].
The $\operatorname{Rep}(\mathbf{T})$-action employed in the Tannakian formalism of [Bez16] described above corresponds to the action of the commutative subalgebra generated by $\theta_{i}^{\pm 1}$. The action of $V\in\operatorname{Rep}(\mathbf{G})$ corresponds to the action of the central element in $H^{\textrm{aff}}_{n}$ corresponding to $V$ – see [Lus83], [Gai01]. This element can be expressed as a symmetric Laurent polynomial in $\theta_{i}$, according to the character of $V$, which is expressed on the categorical level as the compatibility condition on $\operatorname{Rep}(\mathbf{T})$- and $\operatorname{Rep}(\mathbf{G})$- actions mentioned above.
Remark 1.3.1.
To speak of a categorification of the homomorphism $\varpi_{alg}$ in the setting of geometric Hecke categories, one needs to consider “mixed” versions of these categories. We do not treat the “mixed” categorification in the present article. We expect, however, that methods of [HL23] can be used to obtain “mixed” versions of our results.
@@
The action of $\operatorname{Rep}(\mathbf{T})$ is defined in Section 2.2 using the description of the homomorphism $\varpi_{\textrm{br}}$ and the braid group action on $\mathcal{H}_{n}$ defined by Rouquier [Rou06]. To any braid $\beta\in B_{n}$ one can attach an object $F_{\beta}\in\mathcal{H}_{n}$ such that the action of $\beta\in B_{n}$ is isomorphic to the functor of convolution with $F_{\beta}$. For any weight $\boldsymbol{\lambda}$} there is a lift $\tilde{\theta}({\boldsymbol{\lambda}})\in B_{n}^{\textrm{aff}}$ of the Bernstein commutative subalgebra element $\theta({\boldsymbol{\lambda}})\in H_{n}^{\textrm{aff}}$ attached to $\boldsymbol{\lambda}$}. We define the action of $\boldsymbol{\lambda}$}, considered as an object in $\operatorname{Rep}(\mathbf{T})$, to be the functor corresponding to the action of the braid $\varpi_{\textrm{br}}(\tilde{\theta}({\boldsymbol{\lambda}}))$. It is isomorphic to the functor of convolution with $F_{\varpi_{\textrm{br}}(\tilde{\theta}({\boldsymbol{\lambda}}))}$. This complex can be expressed as a convolution of complexes of the form $F_{\beta}$ attached to the Jucys–Murphy braids.
@@
The operation $\star$ is sometimes called $\ast$-convolution, to distinguish it from the dual operation using $m_{!}$ instead of $m_{*}$ in its definition. In this paper, we will work with the $\ast$-convolution -$\star$- and refer to it simply as convolution. This operation is a categorification of the operation of convolution of functions on $G$.
The equivariant category $D^{b}_{G}(G)$, where $G$ acts on itself by the adjoint action, can be equipped with the compatible monoidal structure. The resulting monoidal category is canonically braided. In particular, for any pair of objects $\mathcal{F},\mathcal{G}\in D^{b}_{G}(G)$, there is a braiding isomorphism
@@
$\theta:\operatorname{Id}_{D^{b}_{G}(G)}\to\operatorname{Id}_{D^{b}_{G}(G)},$
@@
See [BD14] for a detailed discussion.
The convolution $\star$ on $D^{b}(G)$ is right $t$-exact with respect to the perverse $t$-structure, since $m$ is an affine morphism. This allows us to equip $P(G),P_{G}(G)$ with a perversely truncated convolution
@@
This makes $P_{G}(G)$ a braided monoidal category as well. The canonical braided structure on $D^{b}_{G}(G)$ is almost never symmetric, even in the case when $G$ is commutative.
It is well known that $G=\textrm{GL}_{n}$ satisfies the following property: for any $g\in G$, the centralizer $Z_{G}(g)$ is connected. We check in Section 3.2 that in this case $P_{G}(G)$ with the monoidal structure $\star^{0}$ is symmetric braided. This essentially follows from the fact that the twist automorphism $\theta$ acts by the identity on perverse objects. Consequently, we can speak of various Schur functors in this category, in particular of exterior powers $\wedge^{k}\mathcal{F}$ for any object $\mathcal{F}\in P_{G}(G)$. In order to construct the action of $\operatorname{Rep}(\mathbf{G})$, we compute the exterior powers of a particular object called the parabolic Springer sheaf (with respect to a particular parabolic subgroup).
@@
Let $\pi:\tilde{\mathcal{N}}_{P}\to G$ be the natural projection, which is known to be a semismall map, and define a perverse sheaf $\mathcal{S}_{P}=\pi_{*}\underline{\overline{\mathbf{Q}}_{\ell}}_{\tilde{\mathcal{N}}_{P}}[2\dim U].$ Let $\mathcal{N}_{G}$ stand for the variety of unipotent elements in $G$. The sheaves $\mathcal{S}_{P}$ are supported on $\mathcal{N}_{G}.$
Now let $V=\mathbf{k}^{n}$. The Weyl group $W\simeq S_{n}$ acts on $V$ by permuting the standard basis. Let $W_{1}\subset W$ be the stabilizer of the first basis vector. Let $G=\textrm{GL}_{n}=\textrm{GL}(V)$, let $B$ be the Borel subgroup stabilizing the standard complete flag in $V$, and let $P_{1}\supset B$ be the parabolic subgroup fixing the line in $\mathbb{P}(V)$ spanned by the first basis vector. Write
@@
Let
$\operatorname{Rep}(W)\to D^{b}_{G}(\mathcal{N}_{G}),V^{\prime}\mapsto\mathcal{S}_{V^{\prime}}$
stand for the functor given by the Springer correspondence [Bry86], [BM81]. We follow the conventions of [Bry86], so that the trivial representation of $W$ corresponds to the punctual perverse sheaf at the unit $e\in\mathcal{N}_{G}$. Since $V\simeq\operatorname{Ind}_{W_{1}}^{W}\operatorname{triv}_{W_{1}},$ we have $\mathcal{S}_{n}\simeq\mathcal{S}_{V}$ [BM81]. Write $\mathcal{S}_{\wedge^{k}}:=\mathcal{S}_{\wedge^{k}V}$.
We will look for a $\operatorname{Rep}(\mathbf{G})$-action on $\mathcal{H}_{n}^{\mathrm{perf}}$ such that the convolution with $\mathcal{S}_{n}$ gives the action of the defining representation $\mathbf{V}$ of $\mathbf{G}=\mathbf{GL}(\mathbf{V})$. Our main computational result is the following
@@
Exterior powers of $\mathcal{S}_{n}$ with respect to the symmetric monoidal structure $\star^{0}$ satisfy the following: for $k\geq 0,$ there is an isomorphism
$\wedge^{k}\mathcal{S}_{n}\simeq\mathcal{S}_{\wedge^{k}}.$
It is well-known that the category $D^{b}_{G}(G)$ acts on $\mathcal{H}_{n}$ by left and right convolution. Moreover, this action is central, meaning that the action on the right can be identified with the action on the left, using the equivariance with respect to adjoint action. It is thus natural to look for a $\operatorname{Rep}(\mathbf{G})$-action on $\mathcal{H}_{n}$ that factors through the action of $D^{b}_{G}(G)$.
Let $\operatorname{Rep}(\mathbf{G})^{+}$ stand for the symmetric monoidal category of polynomial representations of $\mathbf{G}=\mathbf{GL}_{n}$. We have the following corollary of Theorem 1.3.2 (Section 3.5.1 and the proof of Corollary 3.5.2).
@@
There is a symmetric monoidal functor $\operatorname{Rep}(\mathbf{G})^{+}\to P_{G}(G)$ sending the defining representation $\mathbf{V}$ of $\mathbf{G}$ to $\mathcal{S}_{n}$.
@@
Remark 1.3.4.
It would be interesting to consider the values of other Schur functors applied to $\mathcal{S}_{n}$. The author, however, does not know how to approach their computation, even in the “linearized” case of sheaves on the Lie algebra $\mathfrak{g}$, as in [BT24]. In that case, for example, the author does not expect other Schur functors in the derived category to be perverse and does not see a reason to expect them to be semisimple.
Remark 1.3.5.
We note that the functor $\varpi$ constructed with the help of the $\operatorname{Rep}(\mathbf{G})$-action above turns out to have a very simple description, see Section 4.5. The non-trivial part of the present paper is the construction of the $\operatorname{Rep}(\mathbf{G})$-action via the exterior powers computation, and its various compatibilities.
1.3.3  Jucys–Murphy filtrations and monodromy.
To complete the construction of the linear structure and the functor $\varpi$ one needs to establish the compatibility of $\operatorname{Rep}(\mathbf{G})$ and $\operatorname{Rep}(\mathbf{T})$ actions. This is done in Section 4. The main idea is to use the fact that the objects $\wedge^{k}\mathcal{S}_{n}$ come from the parabolic induction from the Levi subgroup $\textrm{GL}_{k}\times\textrm{GL}_{n-k}$ of $G$. This allows us to construct filtrations of the functors defining the $\operatorname{Rep}(\mathbf{G})$-action by the functors defining the $\operatorname{Rep}(\mathbf{T})$-action categorifying the weight filtration. The monodromy transformations required by the Tannakian formalism of [Bez16] come from the monodromy along the central $\mathbb{G}_{m}\subset\textrm{GL}_{k}$.